OpenStudy (honeymonster):

please help..again..find the value of x for this linear equation...posted in replies

OpenStudy (honeymonster):

\[\frac{ 1 }{ 2 } x + 1=3\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Where is the equation? Please post!

OpenStudy (honeymonster):

it is posted?

OpenStudy (unklerhaukus):

take one from both sides of the equation, then multiply both sides by 2

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

Unkl. may be it would be easier to use the actual transfer ?

OpenStudy (anonymous):

x/2 + 1 = 7/2 => x+2/2 =7/2 => x+2 = 7 => x=7-2 => x=5

OpenStudy (anonymous):

Didn't I just do that?

OpenStudy (cruffo):

Never mind, looks like the /2 on x just went missing :)

OpenStudy (anonymous):

I am sorry, my bad. It's (x+2)/2. :) And did you get it @honeymonster ?

OpenStudy (honeymonster):

err...i got x=5 ? :S

OpenStudy (anonymous):

Yeah, that is correct!

OpenStudy (honeymonster):

ok thank youu :)

OpenStudy (anonymous):

Pleasure :)